• (Problem 1) Given the Lotka-Voltera model: = F(a-cs) dF dt ds dt = S(-k + \F). (1) Linearize the model about the equilibriuDo part (1) and (3)
Answer #1

df dt = F (a-cs) ds = S(-k+7F) dt 1) (FS) =(0,0) Linearize it a sound 10,0) dF dt af as - KS dt 3) If we increase k so popula